OpenStudy (anonymous):

A square plywood platform has a perimeter that is 6 times the length of a side, decreased by 6. Find the length of a side. (Recall the formula P = 4s.)

OpenStudy (anonymous):

initial perimeter will be p1=4s when side is s now side has decreased to (s-6) now perimeter isp2= 6s so equate both the perimeters \[p _{_{1}}=p _{2}\] \[4s=6(s=6)\] \[4s=6s-36\] \[2s=36\] \[s=18\] so the side will be 18.